OpenStudy (anonymous):

PLZ HELP ME!!!!!!!!!!!! If (6y-11)(6y+11)=ay^2-b, what is the value of a?

OpenStudy (isaidavila):

Do you remember that property: (a+b)(a-b)=a^2-b^2

OpenStudy (anonymous):

Sort of nit really

OpenStudy (abdullahm):

Hint: (a-b)(a+b) = a^2 - b^2

OpenStudy (sven101):

ah you beat me to it!

OpenStudy (anonymous):

*not

OpenStudy (anonymous):

halp meeeeee

OpenStudy (isaidavila):

Just apply it to the exercise

OpenStudy (anonymous):

Ohhhh Is it 6?

OpenStudy (anonymous):

Or 6y?

OpenStudy (isaidavila):

It would be the square of 6

OpenStudy (abdullahm):

So, basically to solve your question, you have 6y as a. And then you have (6y)^2 = ay^2 So what is a?

OpenStudy (abdullahm):

(6y)^2 = ay^2 Solve forma.

OpenStudy (anonymous):

-85????

OpenStudy (abdullahm):

Solve for a*

OpenStudy (abdullahm):

@mathgirl_123 what did you do to get that number? It's wrong btw.

OpenStudy (sven101):

mathway.com

OpenStudy (anonymous):

I'm getting 6y idk what I'm doing wrong

OpenStudy (abdullahm):

Ok, ok. What is (6y)^2 = ??

OpenStudy (anonymous):

36y

OpenStudy (abdullahm):

Also, you're not telling us what you are doing so we don't know what you're doing wrong.

OpenStudy (abdullahm):

No, (6y)^2 = 36y^2 So we have 36y^2 = ay^2 Can you tell me what a is?

OpenStudy (anonymous):

(6y)^2=36y

OpenStudy (abdullahm):

No, it doesn't. (ab)^2 =a^2b^2

OpenStudy (anonymous):

Don't you divide 36y^2 by a which is 6y?

OpenStudy (anonymous):

.

OpenStudy (abdullahm):

No... 36y^2 = ay^2 What happens when you divide both sides by y^2? What are you left with?

OpenStudy (anonymous):

I am so sorry I'm having like a blank brain moment I am so confused

OpenStudy (anonymous):

I bet the answer is really obvious idk why I can't figure it out

OpenStudy (anonymous):

.

OpenStudy (anonymous):

Is it 36?

OpenStudy (abdullahm):

It is. :)